OpenStudy (anonymous):

Which of the following represents the area common to the two regions bounded by r = -6 cos(x) and r = 2 - 2 cos(x)?

OpenStudy (dumbcow):

 here is a graph of both functions from 0 to 2pi to find where they intersect just set them equal to each other -6cos = 2-2cos -4cos = 2 cos = -1/2 --> x = 2pi/3 and 4pi/3